# The Sunburst Arc with JWST: III. An Abundance of Direct Chemical Abundances

### v2

@@ 4.1.2 Comparison of Electron Temperatures & Densities
One notable exception to this trend is K, which is an outlier in our measurements. At high densities (), dielectronic recombination contributes substantially to the [OII] emission (Liu et al., 2001); this artificially increases the inferred O+ electron temperature. All of our densities are above this threshold, thus we conclude that the O+ temperature is not reliable for this object.
We also have a measurement of the Balmer jump, which is sensitive to the electron temperature in the hydrogen gas across the whole nebula. We find that the Balmer temperature is closer to the measurements from low ionization gas, though the larger uncertainty on the Balmer temperature means that it is still consistent with the high ionization temperatures within uncertainty.
@@ 4.2.2 Nitrogen Abundance
Because we do not detect the [Nii] line, we use the temperature from the [Sii] ratio as our low-ionization temperature when calculating the nitrogen abundance. One potential caveat when using the S+ temperature is that the lower bound of the S+ ionization energy (10.36 eV) is lower than the hydrogen ionization energy (13.6 eV). Thus S+ can form outside the H+ boundary. We note however that the Balmer jump, sensitive to the electron temperature of H+, yields a temperature that is consistent within uncertainties of the S+ temperature. Thus we would not expect a significant change in our measured N+ abundance from this effect.
@@ 4.2.3 Sulfur Abundance
Our spectrum contains the temperature sensitive lines for both S+ and S++, which we use to calculate the abundance of the lower ionization states of sulfur. We do not have a measurement of a higher ionization state of sulfur, so we must correct for the unobserved high ionization sulfur. There may be a small contribution from S+3 ( eV ionization energy). We use the ionization correction factor (ICF) of Izotov et al. (2006), specifically the “low-Z” component of their Equation 20, to make this correction. We find that the ICF is consistent with 1 within uncertainties, indicating that we have measured the primary states of sulfur within the nebula.
We use the direct S+ and S++ temperatures to measure the abundances of these sulfur ions.
@@ 4.2.4 Argon Abundance
We detect both [Ariii] and [Ariv], allowing us to calculate the ionic abundances of both Ar++ and Ar+3. We correct for unobserved higher ionization states of argon using the ICF in Equation 23 of Izotov et al. (2006), specifically using the “low-Z” component based on our measured oxygen abundance. Most of the gas-phase argon is in the two ionization states which we have measured, resulting in an ICF of and a decreased systematic uncertainty on the total argon abundance.
@@ 4.2.5 Neon Abundance
We only detect neon in the Ne++ state, requiring a correction for unobserved ionization states. We again use the ICF of Izotov et al. (2006) for this correction. The neon abundance and ICF are reported in Table 3. We use the electron temperature from O++ to calculate the neon ionic abundance. While we have a direct measurement of , this temperature has a relatively large uncertainty (owing to the faintness of the [Neiii] line). Additionally, it is consistent within uncertainties of . Thus using the doubly ionized oxygen temperature reduces the statistical uncertainty on the neon abundance measurement without introducing any additional systematic uncertainty, as the two temperatures have been found to be consistent.
@@ 4.2.6 Iron Abundance
We calculate the gas-phase iron abundance based on the strength of the nebular [Feiii] line. To account for the unobserved ionization states of iron, we use the ICF from Equation 24 of Izotov et al. (2006). The ionization energy of Fe++ (16.2–30.7 eV) overlaps both the low and intermediate ionization zones. We choose to use the low-ionization temperature from S+ for the iron abundance calculation, though the final result would not change significantly if we instead used the S++ temperature.
@@ 5 Discussion
5.1 High-Quality Abundances of Multiple Species at
We have detected the auroral lines [Sii], [Oiii], [Siii], and [Oii], allowing us to measure the electron temperature in the low, intermediate, and high ionization zones of the nebula. The resulting temperatures are reported in Table 3, and plotted in Figure 3 alongside temperature measurements from local HII regions from the CHAOS project (Berg et al., 2015; Croxall et al., 2015, 2016; Berg et al., 2020). We do not detect [Nii], so we only set an upper limit on the N+ temperature.
@@ 5.1 High-Quality Abundances of Multiple Species at
This combination of directly measured electron temperatures for each of the low, intermediate, and high ionization zones, plus detection of the relevant emission lines for the primary states of multiple elements, enables the same quality of abundance measurements as has been used in nearby galaxies. We can therefore directly compare the gas-phase abundances of this young star cluster at to abundances measured in nearby HII regions. This is the first time chemical abundance patterns have been measured to this level of precision in a galaxy beyond the local universe.
We plot the abundances of nitrogen, sulfur, argon, neon, and iron relative to oxygen in Figure 4. We compare our abundances to measurements from local HII regions from CHAOS (Berg et al., 2015; Croxall et al., 2015, 2016; Berg et al., 2020), low-redshift galaxies from SDSS (Izotov et al., 2006), and other galaxies measured at cosmic noon (Sanders et al., 2023b; Rogers et al., 2024; Welch et al., 2024). Solar abundances from Asplund et al. (2021) are also marked. We find that our measured abundances of sulfur, argon, and neon are all consistent with solar abundance patterns. Additionally, these elements are consistent with local HII regions, indicating the enrichment of alpha elements in the LCE is following standard patterns, in contrast with the low alpha abundances for a galaxy reported in Rogers et al. (2024). Our nitrogen abundance is high compared to local HII regions, similar to the galaxies reported in Sanders et al. (2023b). We discuss this enhanced nitrogen abundance in the following section. We find an iron abundance that is slightly below the solar value, however it matches the Fe/O abundances of low-redshift galaxies with similar oxygen abundances (Izotov et al., 2006). Our measured Fe/O is still higher than what has been observed in the Magellanic Clouds (Domínguez-Guzmán et al., 2022), however our measured dust is also lower, indicating that there is less dust available onto which the iron can condense.
@@ 5.2 Nitrogen Enrichment from Wolf-Rayet Stars
The oxygen abundance of the LCE places it just below the point at which measurements of local HII regions have begun to find significant contribution from secondary nitrogen production. We might therefore expect it to only be affected by primary nitrogen production, placing it along the N/O plateau for low-metallicity galaxies and HII regions (e.g. within uncertainties of the photoionization model prediction from Nicholls et al. (2017)). Additionally, the young stellar age that has been previously reported for this star cluster ( Myr, Chisholm et al. (2019)) would typically be indicative of only primary nitrogen production driving the N/O abundance. The elevated N/O measured for the LCE is in tension with this interpretation, indicating another source of nitrogen enrichment is driving the abundance.
@@
Elevated nitrogen abundances have been observed previously in low-redshift galaxies (e.g., Brinchmann et al., 2008; Amorín et al., 2010; Berg et al., 2011), galaxies at cosmic noon (e.g., Masters et al., 2014; Sanders et al., 2023b), and more recently in high redshift galaxies (Cameron et al., 2023; Senchyna et al., 2023; Marques-Chaves et al., 2024; Castellano et al., 2024). The elevated nitrogen has often been attributed to nitrogen enrichment from Wolf-Rayet stars (Henry et al., 2000; Kobayashi & Ferrara, 2024). However it has also been suggested that inflows of low-metallicity gas can dilute the oxygen abundance, driving increases in the N/O ratio without requiring enrichment from Wolf-Rayet stars (e.g., Köppen & Hensler, 2005; Amorín et al., 2010; Andrews & Martini, 2013).
@@
A companion paper utilizing the same stacked spectrum determined that the LCE contains spectral signatures of Wolf-Rayet stars (Rivera-Thorsen et al., 2024). Previous analyses of galaxies showing WR signatures at low redshift have found that abundances of alpha elements (Ar & Ne) relative to oxygen were not elevated (Brinchmann et al., 2008). Recent theoretical analysis of WR abundance patterns from Kobayashi & Ferrara (2024) find a similar result, that only the nitrogen abundance will change significantly with enrichment from WR stars. We see a similar abundance pattern here, with only N/O elevated relative to similarly low-O/H galaxies and HII regions. The similar abundance pattern could be indicative of WR stars driving the elevated N/O in the Sunburst LCE. However earlier studies have suggested that the WR stellar winds driving the enhanced nitrogen abundance would also carry significant helium (Kobulnicky et al., 1997). We find a helium abundance that is consistent with solar (, Asplund et al., 2021). Brinchmann et al. (2008) also find no evidence for elevated helium abundance in their WR galaxy sample, yet still conclude that the elevated N/O is likely due to WR stellar winds. We thus make the same conclusion here, that WR stellar winds are driving increased N/O abundance in the Sunburst LCE. A recent analysis of several high-redshift () galaxies with elevated N/O also found evidence for elevated He/H (Yanagisawa et al., 2024). These authors suggest that elevated densities () could be related to the high helium abundance. We measure a similar density () but see no increase in helium abundance.
@@
A previous analysis of the LCE using rest-frame UV spectra also measured a high N/O abundance (Pascale et al., 2023). However, their measurement () is higher than our direct measurement by dex, though the two are still consistent within . Pascale et al. (2023) hypothesize that the nitrogen enrichment is most enhanced in the high-density, high-ionization regions of the cluster. Our measurement using N+ being slightly lower than their measurement based on N++ would be consistent with this picture. However we also note that the [NII] transitions used here have lower critical densities than the higher-ionization lines, so collisional de-excitation could be lowering the measured flux. Given the high measured densities here, it is possible that our measured N/O is slightly low, owing to the effects of collisional de-excitation.

### v3

@@ 4.1.2 Comparison of Electron Temperatures & Densities
The differences in temperature between high-ionization and low- and intermediate-ionization gas are shown in Figure 3. The low and intermediate ionization states of sulfur have consistent temperatures, generally matching the trend seen in nearby Hii regions from the CHAOS sample (Berg et al., 2020). The upper limit placed on the N temperature implies that the Sunburst LCE does not fit neatly with the local N+ vs S++ correlation, as the actual N+ temperature would be low relative to . However, Berg et al. (2020) find increased scatter in this relation in high-ionization nebulae (determined by increased ), which is consistent with our finding for the highly ionized Sunburst LCE. Finally, we see that the high-ionization temperature of O++ is significantly elevated relative to the local O++ vs S++ relation (middle panel of Figure 3). The local Hii regions from Berg et al. (2020) do not include regions as highly ionized as the Sunburst LCE (their reported flux ratios only reach , while the Sunburst LCE ratio is nearly an order of magnitude higher). In higher ionization galaxies, Berg et al. (2021) find larger temperature gradients between the higher and lower ionization zones, similar to what we observe. Additionally, we find that the Ne++ temperature is consistent with the O++ temperature, corroborating that the high ionization gas is truly hotter than the lower ionization zones. We therefore conclude that the strong radiation sources within the stellar cluster powering the Sunburst LCE are driving the high ionization gas near the center of the nebula to higher temperatures than the lower-ionization gas towards the edge of the region.
One notable outlier in our measurements is K. At high densities (), dielectronic recombination contributes substantially to the [OII] emission (Liu et al., 2001); this artificially increases the inferred O+ electron temperature. All of our densities are above this threshold, thus we conclude that the O+ temperature is not reliable for this object. However, we note that the recombination contribution would have to be substantial to account for the full offset between our measured ) and the other low- and intermediate-ionization zone temperatures from S. It is therefore likely that other effects are contributing the this offset. One potential contributor is the nebular dust attenuation, since the wavelength gap between [Oii] and [Oii] is large. Our measurement of dust attenuation assumes a Cardelli et al. (1989) dust law, however recently Sanders et al. (2024) found that the nebular attenuation curve for a galaxy differed substantially from locally calibrated models. An altered attenuation curve would potentially bias the O+ temperature that we measure. However such a dust attenuation curve measurement is beyond the scope of the current work.
The hydrogen gas temperature measured from the Balmer jump is lower than our other electron temperature measurements. The Balmer temperature is nearest to the other well-measured low-ionization temperature from S+ ( K); these two temperatures are consistent within . The upper limit on K is also similar to the Balmer temperature, though without a measured [Nii] line the N+ temperature remains uncertain. The Balmer temperature is significantly lower than the measured high-ionization temperature from O++. Interestingly, Hayes et al. (2024) find a similar discrepancy between the high-ionization temperature and the Bamer jump temperature. However, their stacked galaxy spectra show an anticorrelation between stellar population age and Balmer jump temperature, with which our measurement would appear inconsistent. Lastly, we note that Guseva et al. (2006) find no significant discrepancies between temperatures measured from the Balmer jump and those from O++. It is possible that the 55 spaxel apertures used to create the spectrum analyzed here is large enough that the gas in the outskirts of the nebula biases the measured temperature low. Varying the size of the extraction aperture could test this hypothesis, however we leave this test for future work.
@@ 4.2.2 Nitrogen Abundance
Because we do not detect the [Nii] line, we use the temperature from the [Sii] ratio as our low-ionization temperature when calculating the nitrogen abundance. One potential caveat when using the S+ temperature is that the lower bound of the S+ ionization energy (10.36 eV) is lower than the hydrogen ionization energy (13.6 eV). Thus S+ can form outside the H+ boundary. We note however that the Balmer jump, sensitive to the electron temperature of H+, yields a temperature that is consistent within of the S+ temperature. Thus we would not expect a significant change in our measured N+ abundance from this effect. Additionally, we calculated the N abundance using the upper limit on K, and find that this is consistent with the abundance measured using .
@@ 4.2.3 Sulfur Abundance
Our spectrum contains the temperature sensitive lines for both S+ and S++, which we use to calculate the abundance of the lower ionization states of sulfur. We use the direct S+ and S++ temperatures to measure the abundances of these sulfur ions.
We do not have a measurement of a higher ionization state of sulfur, so we must correct for the unobserved high ionization sulfur. There may be a small contribution from S+3 ( eV ionization energy). We use the ionization correction factor (ICF) of Izotov et al. (2006), specifically the “low-Z” component of their Equation 20, to make this correction. We find that the ICF is consistent with 1 within uncertainties. As a consistency check, we calculate the sulfur ICF using the corrections reported in Amayo et al. (2021). We find that the ICF is still consistent with 1 within uncertainties.
These ICFs indicate that the majority of the sulfur in this Hii region exists in the low and intermediate ionization states. However the presence of higher ionization [Ariv] suggests we might expect a more significant contribution from the higher ionization S+3 state. It is possible that the previously described ICFs are insufficient to describe the high ionization of the Sunburst LCE. To further explore this, we consider the 4-zone photoionization models of Berg et al. (2021). Using their [Ariv]/[Ariii] indicator for in the higher ionization region gives for the Sunburst LCE, which would indicate an ICF(S) . This higher ICF yields , slightly more than above our estimate based on the ICF of Izotov et al. (2006) reported in Table 3. A more specific photoionization model for this source would likely be needed to more accurately constrain the contribution of higher ionization states of sulfur, however for our present analysis we conclude that any potential offset is likely contained within our measurement uncertainties.
@@ 4.2.4 Argon Abundance
We detect both [Ariii] and [Ariv], allowing us to calculate the ionic abundances of both Ar++ and Ar+3. We correct for unobserved higher ionization states of argon using the ICF in Equation 23 of Izotov et al. (2006), specifically using the “low-Z” component based on our measured oxygen abundance. Most of the gas-phase argon is in the two ionization states which we have measured, resulting in an ICF of and a decreased systematic uncertainty on the total argon abundance. The ICF of Amayo et al. (2021) provides a similar conclusion that we are observing most of the Ar in the nebula, as the ICF is again consistent with 1 within uncertainties.
@@ 4.2.5 Neon Abundance
We only detect neon in the Ne++ state, requiring a correction for unobserved ionization states. We again use the ICF of Izotov et al. (2006) for this correction, and check for consistency using the ICF of Amayo et al. (2021). The neon abundance and ICFs are reported in Table 3. The two ICF estimates are consistent within uncertainties. We use the electron temperature from O++ to calculate the neon ionic abundance. While we have a direct measurement of , this temperature has a relatively large uncertainty (owing to the faintness of the [Neiii] line). Additionally, it is consistent within uncertainties of . Thus using the doubly ionized oxygen temperature reduces the statistical uncertainty on the neon abundance measurement without introducing any additional systematic uncertainty, as the two temperatures have been found to be consistent.
@@ 4.2.6 Iron Abundance
We calculate the gas-phase iron abundance based on the strength of the nebular [Feiii] line. To account for the unobserved ionization states of iron, we use the ICF from Equation 24 of Izotov et al. (2006). No Fe ICF is reported in Amayo et al. (2021). The ionization energy of Fe++ (16.2–30.7 eV) overlaps both the low and intermediate ionization zones. We choose to use the low-ionization temperature from S+ for the iron abundance calculation, though the final result would not change significantly if we instead used the S++ temperature.
@@
Berg et al. (2021) calculate Fe ICFs as a function of ionization parameter using photoionization models of two nearby extreme emission line galaxies. Their ICFs have a strong dependence on ionization parameter when only Fe++ is observed, as is the case for the Sunburst LCE. Using our estimate of based on the [Ariv]/[Ariii] indicator from Berg et al. (2021) along with their Fe ICF, we find that the correction could be as high as . We therefore encourage caution be used in interpreting our reported iron abundance.
@@ 4.3 Helium Abundance
We have detected the auroral lines [Sii], [Oiii], [Siii], and [Oii], allowing us to measure the electron temperature in the low, intermediate, and high ionization zones of the individual Hii region of the Sunburst LCE. The resulting temperatures are reported in Table 3, and plotted in Figure 3 alongside temperature measurements from local Hii regions from the CHAOS project (Berg et al., 2015; Croxall et al., 2015, 2016; Berg et al., 2020). We do not detect [Nii], so we set a upper limit on the N+ temperature. As discussed in Section 4.1.2, we see a gradient in temperatures between the high-ionization regions at the center of the nebula and the lower-ionization regions towards the edge of the nebula. This gradient is likely driven by the strong radiation produced by the population of massive stars in the central star cluster in this region.
@@
This combination of directly measured electron temperatures for each of the low, intermediate, and high ionization zones, plus detection of the relevant emission lines for the primary states of multiple elements, enables the same quality of abundance measurements as has been used in nearby galaxies. We can therefore directly compare the gas-phase abundances of the Hii region surrounding this young star cluster at to abundances measured in nearby Hii regions. This is the first time chemical abundance patterns have been measured to this level of precision in an individual Hii region beyond the local universe.
We plot the abundances of nitrogen, sulfur, argon, neon, and iron relative to oxygen in Figure 4. We compare our abundances to measurements from local Hii regions from CHAOS (Berg et al., 2015; Croxall et al., 2015, 2016; Berg et al., 2020), low-redshift galaxies from SDSS (Izotov et al., 2006), and other galaxies measured at cosmic noon (Sanders et al., 2023b; Rogers et al., 2024; Welch et al., 2024). Solar abundances from Asplund et al. (2021) are also marked. We find that our measured abundances of sulfur, argon, and neon are all consistent with solar abundance patterns. Additionally, these elements are consistent with local Hii regions, indicating the enrichment of alpha elements in the LCE is following standard patterns, in contrast with the low Ar/O abundance for a galaxy reported in Rogers et al. (2024). Our nitrogen abundance is high compared to local Hii regions, similar to the galaxies reported in Sanders et al. (2023b). We discuss this enhanced nitrogen abundance in the following section. We find an iron abundance that is slightly below the solar value, however it matches the Fe/O abundances of low-redshift galaxies with similar oxygen abundances (Izotov et al., 2006). Our measured Fe/O is still higher than what has been observed in the Magellanic Clouds (Domínguez-Guzmán et al., 2022), however our measured dust is also lower, indicating that there is less dust available onto which the iron can condense.
@@ 5.1 Nitrogen Enrichment from Wolf-Rayet Stars
The oxygen abundance of the Sunburst LCE places it just below the point at which measurements of local Hii regions have begun to find significant contribution from secondary nitrogen production. We might therefore expect it to only be affected by primary nitrogen production, placing it along the N/O plateau for low-metallicity galaxies and Hii regions (e.g. within uncertainties of the fit to stellar abundance data from Nicholls et al. (2017)). The elevated N/O measured for the LCE is in tension with this interpretation, indicating another source of nitrogen enrichment is driving the abundance.
@@
Elevated nitrogen abundances have been observed previously in low-redshift galaxies (e.g., Brinchmann et al., 2008; Amorín et al., 2010; Berg et al., 2011), galaxies at cosmic noon (e.g., Masters et al., 2014; Sanders et al., 2023b), and more recently in high redshift galaxies (Cameron et al., 2023; Senchyna et al., 2023; Marques-Chaves et al., 2024; Castellano et al., 2024). The elevated nitrogen has often been attributed to nitrogen enrichment from Wolf-Rayet stars (Henry et al., 2000; Kobayashi & Ferrara, 2024), though other populations such as AGB stars (D’Antona et al., 2023) and supermassive stars (Charbonnel et al., 2023) have also been hypothesized as drivers of N enhancement. It has also been suggested that inflows of low-metallicity gas can dilute the oxygen abundance, driving increases in the N/O ratio without requiring enrichment from Wolf-Rayet stars (e.g., Köppen & Hensler, 2005; Amorín et al., 2010; Andrews & Martini, 2013).
@@
A companion paper utilizing the same stacked spectrum determined that the LCE contains spectral signatures of Wolf-Rayet stars (Rivera-Thorsen et al., 2024). Previous analyses of galaxies showing WR signatures at low redshift have found that abundances of alpha elements (Ar & Ne) relative to oxygen were not elevated (Brinchmann et al., 2008). Recent theoretical analysis of WR abundance patterns from Kobayashi & Ferrara (2024) find a similar result, that only the nitrogen abundance will change significantly with enrichment from WR stars. We see a similar abundance pattern here, with only N/O elevated relative to similarly low-O/H galaxies and Hii regions. The similar abundance pattern could be indicative of WR stars driving the elevated N/O in the Sunburst LCE.
Earlier studies have suggested that the WR stellar winds driving the enhanced nitrogen abundance would also carry significant helium (Kobulnicky et al., 1997). We find a helium abundance that is consistent with solar (, Asplund et al., 2021). Brinchmann et al. (2008) also find no evidence for elevated helium abundance in their WR galaxy sample, yet still conclude that the elevated N/O is likely due to WR stellar winds. We thus make the same conclusion here, that WR stellar winds are driving increased N/O abundance in the Sunburst LCE. A recent analysis of several high-redshift () galaxies with elevated N/O also found evidence for elevated He/H (Yanagisawa et al., 2024). These authors suggest that elevated densities () could be related to the high helium abundance. We measure a similar density () but see no increase in helium abundance.
@@
Very massive stars (VMS, ) have been suggested to be core H-burning WN stars (WNh, Vink, 2023). The stellar population fits of Rivera-Thorsen et al. (2024) suggest % of the WR stars in the Sunburst LCE are WNh. Previous analyses have also suggested the Sunburst LCE contains VMS (Meštrić et al., 2023). Estimates of nucleosynthetic yields from VMS suggest that these stars could contribute significantly to the overabundance of N (Vink, 2023). Models of VMS also indicate that these stars produce additional Ne compared to regular massive stars (Higgins et al., 2023). Our measured Ne/O being consistent with solar could be in mild tension with the VMS yields, however we note that Higgins et al. (2023) find strong variation in the Ne production relative to O with changing mass. The IMF could alter the total yields. We therefore cannot rule out contributions from VMS in the Sunburst LCE.
A previous analysis of the LCE using rest-frame UV spectra also measured a high N/O abundance (Pascale et al., 2023). Their measurement () is higher than our direct measurement by dex, though the two are still consistent within . Pascale et al. (2023) hypothesize that the nitrogen enrichment is most enhanced in the high-density, high-ionization regions of the cluster. Our measurement using N+ being slightly lower than their measurement based on N++ would be consistent with this picture. However our measurement is significantly higher than the low-density N/O limit () calculated by Pascale et al. (2023) based on the non-detection of the [Nii] lines.
Our measurements of the abundance pattern of the Sunburst LCE disfavor the supermassive star () and AGB star models that have been suggested to explain elevated nitrogen abundances in high redshift galaxies and proto-globular clusters. The AGB model suggested by D’Antona et al. (2023) requires a longer timescale for nitrogen enrichment ( Myr), which is incompatible with the young stellar population of the Sunburst LCE star cluster ( Myr, Chisholm et al., 2019; Rivera-Thorsen et al., 2024). The supermassive star model better matches the observed age of the central star cluster; the short lifetimes of these supermassive stars would allow them to meaningfully alter the nebular abundances in just a few million years. However the supermassive star enrichment model of Charbonnel et al. (2023) predicts an elevated Ne/O abundance ratio, with the most metal-poor objects showing the largest Ne/O increases due to the strong depletion of O in this model. Our measured Ne/O being consistent with solar appears to be in tension with this model prediction. Additionally, the earlier supermassive star model of Gieles et al. (2018) predicts modest enhancement of He in proto-globular cluster environments with a population of supermassive stars. Our measured He/H being consistent with solar is again in mild tension with this prediction, however we note that the models of Gieles et al. (2018) do allow for enhanced N with very low levels of He enrichment. Thus while our measurements appear to be in tension with the supermassive star model, we cannot conclusively rule it out based on our abundance measurements.